Suppose the European Union has a Nominal GDP of 20 trillion Euros, and their GDP deflator is 125. What is the European Union’s Real GDP?

25 trillion Euros
0.16 trillion Euros
16 trillion Euros

With a Real GDP of $15 trillion in 2015 and of $18 trillion in 2016, the real growth would be ________.

2%
16.7%
20%
Answer #1

- GDP deflator =100* nominal gdp / real gdp

125= 100*20/real gdp

Real gdp= 2000/125= 16

16 trillion euros.

- real growth rate= (18-15)*100/15= 20

20%